I do not understand this at all.
Try solving out system of equation with four variables. x + y + z + w = 0 x + y - z - w = 1; x - y + z + w = 3; x - y - z + w = -1.

Answer by stanbon(75887)

Try solving out system of equation with four variables.
x + y + z + w = 0
x + y - z - w = 1
x - y + z + w = 3
x - y - z + w = -1
------------------------
I used the Matrix function of a TI calculator to get the following:
x = 2
y = -3/2
z = 2
w = -5/2
===================
Cheers,
Stan H.